# On stable assignments generated by choice functions of mixed type

### v1

@@ 4 Rotations
In case (ii) with $w\in V_{L}$, we have $\pi^{\rm c}_{w}(x^{\prime})\geq\pi^{\rm c}_{w}(x)$. Then $e\in\pi^{i}<\pi^{\rm c}_{w}(x)$ implies $e\in\pi^{i}<\pi^{\rm c}_{w}(x^{\prime})$, whence $e$ is not interesting for $w$ under $x^{\prime}$.
In case (iii) with $f\in V_{L}$, we have $e\in H_{f}(x)$ but $e\notin\widetilde{H}_{f}(x)$ (for otherwise there would be $e\in\widetilde{H}_{f}(x)=D_{f}(x)$, implying $x^{\prime}(e)>x(e)$). This implies that $e\in H_{w}(x)$. (Here we appeal to Definition 3 and notice that $w$ cannot be deficit, for otherwise $f$ would be labeled as singular by Cleaning procedure.) Then we are as in case (i).
Finally, in case (iv) with $f\in V_{L}$, compare $\pi^{i}\in\Pi_{f}$ (containing $e$) with $\pi^{\rm c}_{f}(x^{\prime})=\pi^{j}$ (containing $H_{f}(x^{\prime})=D_{f}(x)$). If $\pi^{i}<\pi^{j}$, we are done. So assume that $\pi^{i}\geq\pi^{j}$. The vertex $w$ is not deficit. Also $e\notin D_{f}(x)$ (otherwise we obtain $x^{\prime}(e)>x(e)$). It follows (cf. Definition 3) that the critical tie $\pi^{\rm c}_{w}(x)$ must be better than the tie in $\Pi_{w}$ containing $e$. This implies that $e\in E^{\rm end}_{w}(x^{\prime})$, and $e$ is not interesting for $w$ under $x^{\prime}$.
@@
Let $\lambda:{\cal L}^{\rm rot}(x)\to{\mathbb{R}}_{+}$ satisfy $\lambda(L)\leq\tau_{L}$ for all $L\in{\cal L}^{\rm rot}(x)$. Then $x^{\prime}:=x+\sum(\lambda(L)\Delta_{L}\colon L\in{\cal L}^{\rm rot}(x))$ is a stable assignment.
Proof This is shown by use of Lemma 4.8 and the fact that the maximal strong components of the active graph are pairwise disjoint (the latter implies, in particular, that $x^{\prime}$ does not exceed upper bounds and quotas). More precisely, we have to show that each edge $e=fw\in E$ is non-blocking for $x^{\prime}$. This is immediate if both vertices $f,w$ belong to the same component in ${\cal L}^{\rm rot}$ (by appealing to Lemma 4.8), or if both equalities $x^{\prime}_{f}=x_{f}$ and $x^{\prime}_{w}=x_{w}$ take place (since the stability of $x$ implies that $e$ is not blocking for $x$ and $x^{\prime}$).
@@
For a sufficiently small $\varepsilon>0$, the set $X^{\varepsilon}(x)$ defined in (4.11) consists of stable assignments.
This gives one direction in assertion (a) of Theorem 4.2. The other (easier) direction, saying that any stable assignment $x^{\prime}\in{\cal S}$ close to $x$ and satisfying $x^{\prime}\prec x$ belongs to $X^{\varepsilon}(x)$ for an appropriate $\varepsilon$, can be concluded from our discussion in the beginning of this section (including Lemma 4.1 and Cleaning procedure).
Now we show part (b) of this theorem (which leads to the nice property that any stable $y\prec x$ can be reached from $x$ by a sequence of rotations).
@@
    for $f\in V^{\gtrless}\cap F$, the relation $C_{f}(x_{f}\vee y_{f})=x_{f}$ (in view of $x_{f}\succ y_{f}$) implies $E^{>}_{f}\subseteq H_{f}(x)\cup E^{\rm end}_{f}(x)$ and $E^{<}_{f}\subseteq E^{\rm beg}_{f}(x)\cup\{\pi^{\rm c}_{f}(x)\}$; and
@@
    for $w\in V^{\gtrless}\cap W$, the relation $C_{w}(x_{w}\vee y_{w})=y_{w}$ (in view of $y\succ_{W}x$) implies that either $E^{>}_{w}\subseteq T_{w}(x)$ and $H_{w}(x)\subseteq E^{<}_{w}\subseteq\pi^{\rm c}_{w}(x)$ (when $\pi^{\rm c}_{w}(y)=\pi^{\rm c}_{w}(x)$), or $E^{>}_{w}\subseteq E^{\rm beg}_{w}(x)$ and $E^{<}_{w}=\{e\in\pi^{\rm c}_{w}(x)\colon x(e)>0\}$ (when $\pi^{\rm c}_{w}(y)>\pi^{\rm c}_{w}(x)$).
Note that from there relations one can conclude that
@@
    for any $f\in V^{\gtrless}\cap F$, there holds $E^{>}_{f}\subseteq\widetilde{H}_{f}(x)\cup\cup(\widehat{H}^{i}_{f}(x)\colon\pi^{i}\in\Pi^{\rm end}_{f}(x))$ (where $\widetilde{H}_{f}$ and $\widehat{H}^{i}_{f}$ are defined in Definition 3).
Using these observations, we construct sets $A_{v}$ of active edges for $x$ and $v\in V^{\gtrless}$ by use of the following procedure. Initially, we fix an arbitrary vertex $f=f_{0}$ in $V^{\gtrless}\cap F$. Relation ((4.19)) implies that $D_{f}(x)\neq\emptyset$. Define $A_{f}:=D_{f}(x)$ and $W(A_{f}):=\{w\colon fw\in A_{f}\}$. We assert that
@@
As a result, this procedure constructs a collection ${\cal A}$ of sets $A_{v}$ for certain vertices $v$ in $V^{\gtrless}$. Here if $v=f\in F$, then $A_{f}$ coincides with $D_{f}(x)$, and if $v=w\in W$, then $A_{w}=H_{w}$. Upon termination of the procedure, for each edge $fw\in A_{f}\in{\cal A}$, the collection ${\cal A}$ contains the set $A_{w}$, and similarly, for each edge $fw\in A_{w}\in{\cal A}$, ${\cal A}$ contains $A_{f}$. In other words, ${\cal A}$ determines a subgraph $\overrightarrow{\Gamma}^{\prime}$ of the active graph $\overrightarrow{\Gamma}(x)$, which is closed under the reachability by directed paths. Then $\overrightarrow{\Gamma}^{\prime}$ contains a maximal strong component $L$ of $\overrightarrow{\Gamma}(x)$. We fix $L$, take the rotation $\rho_{L}$ along with a sufficiently small weight $\lambda>0$, and define $x^{\prime}:=x+\lambda\rho_{L}$. We assert that this $x^{\prime}$ is as required.
@@
    for each $w\in V_{L}\cap W$, we have $A_{w}=H_{w}\subseteq E^{<}_{w}$, and $y(e)<x^{\prime}(e)<x(e)$ holds for all $e\in A_{w}$.
In its turn, for each $f\in V_{L}\cap F$, the edges $e\in A_{f}$ satisfy $x(e)<x^{\prime}(e)<b(e)$. We need somewhat stronger relations involving $y$. More precisely, we assert that
@@
    for each $f\in V_{L}\cap F$, any edge $e=fw\in A_{f}$ ($=D_{f}(x)$) with $y(e)<b(e)$ belongs to either $E^{>}\cup H_{w}(y)$ or to a tie in $\Pi^{\rm end}_{w}(y)$.
Indeed, since $e$ is not blocking for $y$, $e$ belongs to at most one tail among $T_{f}(y)$ and $T_{w}(y)$ (cf.((2.6))); therefore, at least one of the following takes place: (i) $e\in H_{f}(y)$; (ii) $e\in H_{w}(y)$; (iii) $e$ belongs to a tie in $\Pi^{\rm end}_{f}(y)$; and (iv) $e$ belongs to a tie in $\Pi^{\rm end}_{w}(y)$.
@@
Let $x,y\in{\cal S}$ and $y\prec x$. Then there is a maximal strong component $L$ in $\overrightarrow{\Gamma}(x)$ and a number $\lambda\leq\tau_{L}$ such that the assignment $x^{\prime}:=x+\lambda\rho_{L}$ satisfies $y\preceq x^{\prime}\prec x$. The maximum possible weight $\lambda$ (subject to $y\preceq x^{\prime}$) is equal to the minimum between the number $\tau_{L}$, the values $(y(e)-x(e))/\rho_{L}(e)$ over $e\in E_{L}$ such that $\rho_{L}(e)>0$ and $y(e)>x(e)$, and the values $(x(e)-y(e))/|\rho_{L}(e)|$ over $e\in E_{L}$ such that $\rho_{L}(e)<0$ and $x(e)>y(e)$. normal-$\Box$
@@
    for $\varepsilon>0$, system (4.5)–(4.8) has a solution;
@@
Indeed, if (i) is valid, then for $y:=x^{\rm max}$, we have $y\prec x$, and Lemma 4.11 implies validity of (ii). Conversely, if system (4.5)–(4.8) is solvable for $\varepsilon>0$, then (by Lemma 4.8) there exists a stable $x^{\prime}$ such that $x^{\prime}\prec x$, yielding (i). The equivalence of (iii) and (iv) can be concluded from Cleaning procedure. The implication (iii)$\to$(ii) follows from Lemma 4.4. The converse implication (ii)$\to$(iii) can be concluded from the discussion in the beginning part of this section.
@@ 5 Additional properties of rotations
We will refer to the transformation $x\rightsquigarrow x^{\prime}:=x+\lambda\rho_{L}$, where $L\in{\cal L}^{\rm rot}(x)$ and $\lambda$ is an admissible (w.r.t. $x$) weight for the rotation $\rho_{L}$, as the rotational shift along $\rho_{L}$ of value $\lambda$. For brevity, we also say that $x^{\prime}$ is obtained from $x$ by the shift by $\lambda\rho_{L}$. If $\lambda=\tau_{L}(x)$, the shift is called full. As before, the sets $F\cap V_{L}$ and $W\cap V_{L}$ are denoted as $F_{L}$ and $W_{L}$, respectively. One can see that
@@
    the maximal admissible weight $\tau_{L}=\tau_{L}(x)$ is equal to the minimum of the following values: (a) $x(e)/|\rho_{L}(e)|$ over all $e\in H_{w}$, $w\in W_{L}$; (b) $(b(e)-x(e))/\rho_{L}(e)$ over all $e\in D_{f}$, $f\in F_{L}$; and (c) $(x(e)-x(e^{\prime}))/(|\rho_{L}(e)|+\rho_{L}(e^{\prime}))$ over all pairs $e,e^{\prime}\in E_{L}$ such that $e\in H_{w}$ and $e^{\prime}\in\pi^{\rm c}_{w}(x)-H_{w}$ among $w\in W_{L}$.
(Here (a) and (b) ensure that $x^{\prime}$ is nonnegative and does not exceed the upper bounds $b$, and (c) provides (when the critical tie preserves) the inclusion $H_{w}(x)\subseteq H_{w}(x^{\prime})$ for $w\in W_{L}$. In other words, in case (c), the increased assignment in any edge of the set $\pi^{\rm c}_{w}-H_{w}$ should not “overlap” the decreased assignments in $H_{w}$.)
Given a stable $x$, the extraction of critical ties and their heads, and, further, the construction of the active graph $\overrightarrow{\Gamma}$, take linear time $O(|E|)$ (assuming that $|V|=O(|E|)$). The task of finding the strong components in $\overrightarrow{\Gamma}$ can also be solved in linear time (the first linear time algorithm for this task is given in [11]). Then we can easily extract the set ${\cal L}^{\rm rot}$ of maximal strong components. For each $L\in{\cal L}^{\rm rot}$, we can efficiently construct the rotation $\rho_{L}$ by applying to system (4.5)–(4.7) a strongly polynomial method in linear algebra (see, e.g., [15]). (At this moment we cannot suggest a more combinatorial efficient method to solve this system.) Finally, using ((5.1)), it is not difficult to compute the maximal rotational weights $\tau_{L}$ for all rotations $\rho_{L}$ in time $O(|E|)$ (taking into account that the sets $V_{L}$ are pairwise disjoint).
As a result, we obtain the following
@@
Given $x\in{\cal S}$, we can perform, in time polynomial in $|E|$, the following tasks: (a) recognize whether $x=x^{\rm max}$ or not, and (b) in case $x\neq x^{\rm max}$, find all rotations $\rho_{L}$, $L\in{\cal L}^{\rm rot}(x)$, and their maximal admissible weights $\tau_{L}$.
Next, based on the analysis from the previous section, one can estimate the number of rotational transformations (shifts) applied in the process of moving from the minimal stable assignment $x^{\rm min}$ to an arbitrary stable $x\in{\cal S}$, in particular, to the maximal assignment $x^{\rm max}$. Here the following observations are of use (cf. ((4.15)) and ((5.1))):
@@
    Let $x^{\prime}$ be obtained from $x\in{\cal S}$ by the shift along the rotation $\rho_{L}$ of value $\lambda$; then:
@@
      if $\lambda=\tau_{L}$, then at least one of the following takes place: (i) for some $f\in F^{=}$, either $\pi^{\rm c}_{f}(x^{\prime})=\pi^{\rm c}_{f}(x)$ and $H_{f}(x^{\prime})\subset H_{f}(x)$, or $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$; (ii) for some $w\in W^{=}$, either $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})\supset H_{w}(x)$, or $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$.
This monotonicity behavior leads to the following:
@@
    in the process of rotational shifts, for each edge $e\in fw\in E$, the updates of the current assignment $x(e)$ can be chronologically divided into two phases: at the first phase, the value at $e$ is monotone non-decreasing (when $e$ occurs in “heads” concerning the part $F$); and at the second phase, the value at $e$ is monotone non-increasing (when $e$ occurs in “heads” concerning the part $W$).
Using ((5.2)) and ((5.3)), we can obtain the following
@@
(i) If $x\in{\cal S}$ is constructed from $x^{\rm min}$ by a sequence of full rotational shifts, then the number of shifts is $O(|E|)$.
(ii) Any $x\in{\cal S}$ can be constructed from $x^{\rm min}$ by $O(|E|)$ rotational shifts.
@@
Proof  To show (i), let us associate with a full shift in a given sequence one edge $e=fw$ along with some of the following events: 1) $e$ is inserted initially in the head $H_{f}$, where $f\in F$; 2) $e$ leaves the head $H_{f}$; 3) $e$ is inserted initially in the head $H_{w}$, where $w\in W$; and 4) $e$ leaves the head $H_{w}$. More precisely, if $x^{\prime}$ is obtained from $x$ by a full shift along a rotation $\rho_{L}$, then, considering the corresponding cases in ((5.2))(b), we choose $e\in H_{f}(x)-H_{f}(x^{\prime})$ in case $H_{f}(x^{\prime})\subset H_{f}(x)$; choose $e\in\pi^{\rm c}_{f}(x^{\prime})$ in case $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$; choose $e\in H_{w}(x^{\prime})-H_{w}(x)$ in case $H_{w}(x^{\prime})\supset H_{w}(x)$; and choose $e\in H_{w}(x)$ in case $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$. Then these choices match the events 1)–4), and one can see that each edge is associated with at most four full shifts in the process.

### v2

@@ 4 Rotations
In case (ii) with $w\in V_{L}$, we have $\pi^{\rm c}_{w}(x^{\prime})\geq\pi^{\rm c}_{w}(x)$. Then $e\in E^{\rm end}_{w}(x)$ implies $e\in E^{\rm end}_{w}(x^{\prime})$, whence $e$ is not interesting for $w$ under $x^{\prime}$.
In case (iii) with $f\in V_{L}$, we have $e\in H_{f}(x)$ but $e\notin\widetilde{H}_{f}(x)$ (for otherwise there would be $e\in\widetilde{H}_{f}(x)=D_{f}(x)$, implying $x^{\prime}(e)>x(e)$). This implies that $e\notin T_{w}(x)$, and therefore, $e\in H_{w}(x)$. (Here we appeal to Definition 3 and notice that $w$ cannot be deficit; for otherwise, we would have $D_{f}(x)=\emptyset$, whence $f$ should be inserted in $V^{0}$ by Cleaning procedure.) Then we are as in case (i).
Finally, in case (iv) with $f\in V_{L}$, compare $\pi^{i}\in\Pi_{f}$ (containing $e$) with $\pi^{\rm c}_{f}(x^{\prime})=\pi^{j}$ (containing $H_{f}(x^{\prime})=D_{f}(x)$). If $\pi^{i}<\pi^{j}$, we are done. So assume that $\pi^{i}\geq\pi^{j}$. The vertex $w$ is not deficit. Also $e\notin D_{f}(x)$ (otherwise we obtain $x^{\prime}(e)>x(e)$). It follows (cf. Definition 3) that the critical tie $\pi^{\rm c}_{w}(x)$ must be better than the tie in $\Pi_{w}$ containing $e$, i.e. $e\in E^{\rm end}_{w}(x)$. This implies that $e\in E^{\rm end}_{w}(x^{\prime})$, and $e$ is not interesting for $w$ under $x^{\prime}$.
@@
Let $\lambda:{\cal L}^{\rm rot}(x)\to{\mathbb{R}}_{+}$ satisfy $\lambda(L)\leq\tau_{L}(x)$ for all $L\in{\cal L}^{\rm rot}(x)$. Then $x^{\prime}:=x+\sum(\lambda(L)\rho_{L}\colon L\in{\cal L}^{\rm rot}(x))$ is a stable assignment.
Proof This is shown by use of Lemma 4.10 and the fact that the maximal strong components of the active graph are pairwise disjoint (the latter implies, in particular, that $x^{\prime}$ does not exceed upper bounds and quotas). More precisely, we have to show that each edge $e=fw\in E$ with $x^{\prime}(e)<b(e)$ is non-blocking for $x^{\prime}$. This is immediate if both vertices $f,w$ belong to the same component in ${\cal L}^{\rm rot}$ (by appealing to Lemma 4.10), or if both equalities $x^{\prime}_{f}=x_{f}$ and $x^{\prime}_{w}=x_{w}$ take place (since $e$ is not blocking for $x$, implying that $e$ is not blocking for $x^{\prime}$ as well).
@@
For a sufficiently small $\varepsilon>0$, any assignment $x^{\prime}:=x+\Delta^{\phi,\psi}$ with $(\phi,\psi)\in{\cal P}^{\varepsilon}(x)$ (defined in (4.11)) is stable, i.e. $x^{\prime}\in X^{\varepsilon}(x)$ (defined in (4.1)) .
This gives one direction in assertion (a) of Theorem 4.4. The other (easier) direction, saying that any stable assignment $x^{\prime}\in{\cal S}$ close to $x$ and satisfying $x^{\prime}\prec x$ belongs to $\{x+\Delta^{\phi,\psi}\colon(\phi,\psi)\in{\cal P}^{\varepsilon}(x)\}$ for an appropriate $\varepsilon$, can be concluded from our discussion in the beginning of this section (including Lemmas 4.1 and 4.2 and Cleaning procedure).
Now we show part (b) of this theorem (which will lead to the nice property that any stable $y\prec x$ can be reached from $x$ by a sequence of rotations).
@@
    for $f\in V^{\gtrless}\cap F$, the equality $C_{f}(x_{f}\vee y_{f})=x_{f}$ (in view of $x_{f}\succ y_{f}$) implies $E^{>}_{f}\subseteq H_{f}(x)\cup E^{\rm end}_{f}(x)$ and $E^{<}_{f}\subseteq E^{\rm beg}_{f}(x)\cup\{\pi^{\rm c}_{f}(x)\}$; and
@@
    for $w\in V^{\gtrless}\cap W$, the equality $C_{w}(x_{w}\vee y_{w})=y_{w}$ (in view of $y\succ_{W}x$) implies that either $E^{>}_{w}\subseteq T_{w}(x)$ and $H_{w}(x)\subseteq E^{<}_{w}\subseteq\pi^{\rm c}_{w}(x)$ (when $\pi^{\rm c}_{w}(y)=\pi^{\rm c}_{w}(x)$), or $E^{>}_{w}\subseteq E^{\rm beg}_{w}(x)$ and $E^{<}_{w}=\{e\in\pi^{\rm c}_{w}(x)\colon x(e)>0\}$ (when $\pi^{\rm c}_{w}(y)>\pi^{\rm c}_{w}(x)$).
From there relations one can conclude that
@@
    for any $f\in V^{\gtrless}\cap F$, there holds $E^{>}_{f}\subseteq\cup(D^{i}_{f}(x)\colon\pi^{i}\in\widehat{\Pi}_{f}(x))$.
(Where $D^{i}_{f}(x)$ and $\widehat{\Pi}_{f}(x)$ are defined in Definition 3.) Using these observations, we construct sets $A_{v}$ of active edges for $x$ and $v\in V^{\gtrless}$ by the following procedure. Initially, we fix an arbitrary vertex $f=f_{0}$ in $V^{\gtrless}\cap F$. The inclusion in ((4.20)) implies $D_{f}(x)\neq\emptyset$. Define $A_{f}:=D_{f}(x)$ and $W(A_{f}):=\{w\colon fw\in A_{f}\}$. We assert that
@@
As a result, this procedure constructs a collection ${\cal Q}$ of sets $A_{v}$ for certain vertices $v$ in $V^{\gtrless}$. Here if $v=f\in F$, then $A_{f}$ coincides with $D_{f}(x)$, and if $v=w\in W$, then $A_{w}=H_{w}$. Upon termination of the procedure, for each edge $fw\in A_{f}\in{\cal Q}$, the collection ${\cal Q}$ contains the set $A_{w}$, and similarly, for each edge $fw\in A_{w}\in{\cal Q}$, ${\cal Q}$ contains $A_{f}$. In other words, ${\cal Q}$ determines a subgraph $\overrightarrow{\Gamma}^{\prime}$ of the active graph $\overrightarrow{\Gamma}(x)$, which is closed under the reachability by directed paths. Then $\overrightarrow{\Gamma}^{\prime}$ contains a maximal strong component $L$ of $\overrightarrow{\Gamma}(x)$. We fix $L$, take the rotation $\rho_{L}$ along with a sufficiently small weight $\lambda>0$, and define $x^{\prime}:=x+\lambda\rho_{L}$. We assert that this $x^{\prime}$ is as required.
@@
    each $w\in V_{L}\cap W$ satisfies $A_{w}=H_{w}\subseteq E^{<}_{w}$, and each $e\in A_{w}$ satisfies $y(e)<x^{\prime}(e)<x(e)$.
In turn, for each $f\in V_{L}\cap F$, the edges $e\in A_{f}$ satisfy $x(e)<x^{\prime}(e)<b(e)$. We need somewhat stronger relations involving $y$. More precisely, we assert that
@@
    for each $f\in V_{L}\cap F$, any edge $e=fw\in A_{f}$ ($=D_{f}(x)$) with $y(e)<b(e)$ belongs to either $E^{>}\cup H_{w}(y)$ or $E^{\rm end}_{w}(y)$.
Indeed, since $e$ is not blocking for $y$, $e$ belongs to at most one tail among $T_{f}(y)$ and $T_{w}(y)$ (cf. ((2.6))); therefore, at least one of the following takes place: (i) $e\in H_{f}(y)$; (ii) $e\in H_{w}(y)$; (iii) $e$ belongs to $E^{\rm end}_{f}(y)$; and (iv) $e$ belongs to $E^{\rm end}_{w}(y)$.
@@
Let $x,y\in{\cal S}$ and $y\prec x$. Then there exists a maximal strong component $L$ in $\overrightarrow{\Gamma}(x)$ and a number $\lambda\leq\tau_{L}(x)$ such that the assignment $x^{\prime}:=x+\lambda\rho_{L}$ satisfies $y\preceq x^{\prime}\prec x$. The maximum possible weight $\lambda$ (subject to $y\preceq x^{\prime}$) is equal to the minimum between the number $\tau_{L}(x)$, the values $(y(e)-x(e))/\rho_{L}(e)$ over $e\in E_{L}$ such that $\rho_{L}(e)>0$ and $y(e)>x(e)$, and the values $(x(e)-y(e))/|\rho_{L}(e)|$ over $e\in E_{L}$ such that $\rho_{L}(e)<0$ and $x(e)>y(e)$. $\Box$
@@
    system (4.7)–(4.10) has a solution with $\varepsilon>0$;
@@
Indeed, if (i) is valid, then for $y:=x^{\rm max}$, we have $y\prec x$, and Lemma 4.13 implies validity of (ii). Conversely, if system (4.7)–(4.10) is solvable for $\varepsilon>0$, then (by Lemma 4.10) there exists a stable $x^{\prime}$ such that $x^{\prime}\prec x$, yielding (i). The equivalence of (iii) and (iv) can be concluded from Cleaning procedure and ((4.5)). The implication (iii)$\to$(ii) follows from Lemma 4.6. The converse implication (ii)$\to$(iii) can be concluded from the discussion in the beginning part of this section.
@@ 5 Additional properties of rotations
We will refer to the update $x\rightsquigarrow x^{\prime}:=x+\lambda\rho_{L}$, where $L\in{\cal L}^{\rm rot}(x)$ and $\lambda$ is an admissible weight for $x$ and the rotation $\rho_{L}$, as the rotational shift of weight $\lambda$ along $\rho_{L}$ applied to $x$. We also say that $x^{\prime}$ is obtained from $x$ by the shift by $\lambda\rho_{L}$. If $\lambda=\tau_{L}(x)$, the shift is called full (w.r.t. $x$). As before, we write $F_{L}$ for $F\cap V_{L}$, and $W_{L}$ for $W\cap V_{L}$. One can see that
@@
    the maximal admissible weight $\tau_{L}(x)$ is equal to the minimum of the following values: (a) $x(e)/|\rho_{L}(e)|$ over all $e\in H_{w}$, $w\in W_{L}$; (b) $(b(e)-x(e))/\rho_{L}(e)$ over all $e\in D_{f}$, $f\in F_{L}$; and (c) $(x(e)-x(e^{\prime}))/(|\rho_{L}(e)|+\rho_{L}(e^{\prime}))$ over all pairs $e,e^{\prime}\in E_{L}$ such that $e\in H_{w}$ and $e^{\prime}\in\pi^{\rm c}_{w}(x)-H_{w}$ among $w\in W_{L}$.
(Here (a) and (b) ensure that $x^{\prime}$ is nonnegative and does not exceed the upper bounds $b$, and (c) provides (when the critical tie preserves) the inclusion $H_{w}(x)\subseteq H_{w}(x^{\prime})$ for $w\in W_{L}$. In other words, in case (c), the new (increased) assignment in any edge of $\pi^{\rm c}_{w}-H_{w}$ should not exceed the new (decreased) assignment on an edge of $H_{w}$.)
Given a stable $x$, the extraction of critical ties and their heads, and, further, the construction of the active graph $\overrightarrow{\Gamma}$, take linear time $O(|E|)$. The task of finding the strong components in $\overrightarrow{\Gamma}$ can also be solved in linear time (e.g. using the first linear time algorithm for this task in [14]). Then we can easily extract the set ${\cal L}^{\rm rot}$ of maximal strong components. Next, for each $L\in{\cal L}^{\rm rot}$, we efficiently construct the rotation $\rho_{L}$. To do this, we first solve system (4.7)–(4.10) (in which $F^{+}$ and $W^{+}$ are replaced by $F\cap V_{L}$ and $W\cap V_{L}$, respectively, and one puts $\varepsilon=1$, say). And second, we transform the obtained solution $(\phi,\psi)\in{\mathbb{Q}}_{+}^{F^{+}}\times{\mathbb{Q}}_{+}^{W^{+}}$ into the rotation $\rho_{L}\in{\mathbb{Z}}^{E}$.
(Here the system has size $(n+1)\times n$ (where $n:=|V|$) and integer entries $a_{ij}$ with $|a_{ij}|\leq n$. It is solved by Gaussian eliminations in time $O(n^{3})$ (and all intermediate data during the process are rational numbers with numerators and denominators whose encoding sizes (viz. numbers of bits to record) are of order $O(n^{2}\log n)$, as is shown by Edmonds for a general case, see e.g. [17, Sec. 4.3] or [21]. Now to form the desired rotation, the obtained $(\phi,\psi)$ is scaled to get all entries integer-valued, and then we get rid of the common divisor by use of Euclidean algorithm (which takes $O(n^{3}\log n)$ time in total, cf. [17, Sec. 4.2]). At present we cannot offer a more combinatorial efficient method to construct the rotation related to a strong component $L$.)
Finally, using ((5.1)), one can compute the maximal rotational weights $\tau_{L}$ for all rotations $\rho_{L}$ in time $O(|E|)$ (since the sets $V_{L}$ for $L\in{\cal L}^{\rm rot}$ are pairwise disjoint).
Summarizing the above observations, we obtain the following complexity result.
@@
Given $x\in{\cal S}$, we can fulfill, in time polynomial in $|E|$, the following tasks: (a) recognize whether $x=x^{\rm max}$, and (b) in case $x\neq x^{\rm max}$, find all rotations $\rho_{L}$, $L\in{\cal L}^{\rm rot}(x)$, and their maximal admissible weights $\tau_{L}(x)$. Each entry in a rotation has the encoding size bounded by a polynomial in $|E|$, and similarly for the encoding sizes of intermediate data appeared in the process of computing the rotations.
Next, based on the analysis from the previous section, one can estimate the number of rotational transformations (shifts) applied in the process of moving from the minimal stable assignment $x^{\rm min}$ or from an arbitrary stable $x\in{\cal S}$ to another stable assignment or to $x^{\rm max}$. Here the following observations are of use (cf. ((4.16)) and ((5.1))):
@@
    Let $x^{\prime}$ be obtained from $x\in{\cal S}$ by the shift of weight $\lambda$ along a rotation $\rho_{L}$; then:
@@
      for $\lambda=\tau_{L}(x)$, at least one of the following takes place: (i) for some $f\in F^{=}$, either $\pi^{\rm c}_{f}(x^{\prime})=\pi^{\rm c}_{f}(x)$ and $H_{f}(x^{\prime})\subset H_{f}(x)$, or $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$; (ii) for some $w\in W^{=}$, either $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})\supset H_{w}(x)$, or $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$.
Note that this monotonicity behavior leads to the following:
@@
    in the process of rotational shifts, for each edge $e\in fw\in E$, the updates of current assignments $x(e)$ can be chronologically divided into two phases: at the first phase, the value at $e$ is monotone non-decreasing (when $e$ occurs in “heads” concerning the part $F$); and at the second phase, the value at $e$ is monotone non-increasing (when $e$ occurs in “heads” concerning the part $W$).
Also, using ((5.2)), we obtain the following
@@
(i) If a stable assignment is constructed from another one by a sequence of full rotational shifts, then the number of shifts is at most $2|E|$.
(ii) Any $x\in{\cal S}$ can be constructed from $x^{\rm min}$ by $2|E|$ rotational shifts.
@@
Proof To show (i), consider corresponding cases in ((5.2)) and associate with a full shift $x\rightsquigarrow x^{\prime}$ in a given sequence one edge $e=fw$ as follows. Choose $e\in H_{f}(x)-H_{f}(x^{\prime})$ in case $H_{f}(x^{\prime})\subset H_{f}(x)$; choose $e\in H_{f}(x)$ in case $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$; choose $e\in H_{w}(x^{\prime})-H_{w}(x)$ in case $H_{w}(x^{\prime})\supset H_{w}(x)$; and choose $e\in H_{w}(x^{\prime})$ in case $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$. Then each edge is associated with at most two full shifts in the sequence, whence (i) follows.
To show (ii), it is useful to note that, given $x\in{\cal S}$, the task of constructing $x$ by rotational shifts from $x^{\rm min}$ is equivalent to that of constructing $x^{\rm max}$ from $x$ (for we can “swap” the parts $F$ and $W$ and “reverse” rotations). So we can consider the latter problem. (By the way, if needed, we can efficiently verify whether a function $x$ on $E$ obeying $b,q$ is indeed a stable assignment, by checking the absence of blocking edges.)